OpenStudy (anonymous):

Solve: 4(x − 1) = 2(x + 2) + 6

OpenStudy (anonymous):

x=7 to slow

OpenStudy (anonymous):

\[4x -4=2x +4+6\] - First, we distribute \[4x -4=2x +10\] comebine any like terms on both sides \[4x -2x -4=4\] Move x's to one side, I prefer the left, we subtract 2x from each side \[2x -4=10\] once again, we combine like terms \[2x =14\] we move constant terms to one side, add 4 to each side \[x =7\] we isolate x by dividing each side by 2

OpenStudy (anonymous):

I prefer to show you "how" to solve a problem, not just the answer, which teaches you nothing. But nice work.